OpenStudy (anonymous):

Write the inverse variation equation of the following. Y varies inversely with x and y = 9 when x = 14

OpenStudy (anonymous):

First, the inverse equation: \[x/k=y\] Then, substitute: \[14/k=9\] Divide both sides by 14: \[(14/k)/14=9/14\] \[k=9/14\] ((I hope this is right.....I haven't done inverse since 6th grade so.....))

OpenStudy (dan815):